For the following LP problem, graph the region of feasible solution and solve by the corner-point method.

Maximize z = 5x1 + 8x2
Subject to x1 + x2 ≥ 6
3x1 + 2x2 ≤ 30
2x1 + x2 ≤ 5
x1 , x2 ≥ 0
1
Expert's answer
2020-04-30T20:22:38-0400

"x_1+x_2\geq 6"

"3x_1+2x_2\leq 30"

"2x_1+x_2\leq 5"

"x_1, x_2\geq 0"

Maximize "z=5x_1+8x_2"




Green region shows all solutions of "\begin{cases}
x_1+x_2\geq 6
\\3x_1+2x_2\leq 30
\\
2x_1+x_2\leq 5
\end{cases}"

And red region shows all solutions of "x_1,x_2\geq 0"

We can see that there is no solution of the system "\begin{cases}
x_1+x_2\geq 6
\\3x_1+2x_2\leq 30
\\
2x_1+x_2\leq 5
\\ x_1,x_2\geq 0
\end{cases}" because intersection of these regions is empty.

Therefore, we can’t maximize "z=5x_1+8x_2" because there is no "(x_1,x_2)" , for which all inequalities are satisfied.